DEC 11 OWL Summary

Richard Kessler (rkessler@fnal.gov)
Sat, 11 Dec 1999 06:13:01 -0600 (CST)

Shift Summary for OWL shift, DEC 11, 1999
On Shift: R.Kessler
o COSMICS run started at 5:40am (after wake-up)
o ktev3b disk full as predicted by SKIM team.
o KAMI beam test cables removed. BA7 plugged back into SS.
Next PED run will ask for update because I had changed two
BA2 ped channels in the data base. [recall BA2 is NOT used].
****************